# Constructing stable, high-order finite-difference operators on point clouds over complex geometries

### v1

@@ Abstract
High-order difference operators with the summation-by-parts (SBP) property can be used to build stable discretizations of hyperbolic conservation laws; however, most high-order SBP operators require a conforming, high-order mesh for the domain of interest. To circumvent this requirement, we present an algorithm for building high-order, diagonal-norm, first-derivative SBP operators on point clouds over complex geometries. The algorithm is not mesh-free, since it uses a Cartesian cut-cell mesh to define the sparsity pattern of the operators and to provide intermediate quadrature rules; however, the mesh is generated automatically and can be discarded once the SBP operators have been constructed. Using this temporary mesh, we construct local, cell-based SBP difference operators that are assembled into global SBP operators. We identify conditions for the existence of a positive-definite diagonal mass matrix, and we compute the diagonal norm by solving a sparse system of linear inequalities using an interior-point algorithm. We also describe an artificial dissipation operator that complements the first-derivative operators when solving hyperbolic problems, although the dissipation is not required for stability. The numerical results confirm the conditions under which a diagonal norm exists and study the distribution of the norm’s entries. In addition, the results verify the accuracy and stability of the point-cloud SBP operators using the linear advection equation.
@@ 1 Introduction
The present contribution aims to accelerate both the run and pre-processing time of high-order SBP discretizations. Specifically, we present an algorithm to construct high-order, sparse SBP operators whose nodal degrees of freedom are defined on point clouds. The combination of these two elements — high-order SBP and point-cloud nodes — yields an efficient and stable discretization with considerable flexibility.
@@ 1.1 Related work
In [22], the authors present an algorithm to construct a high-order finite-difference method on point clouds. While they do not use the term summation-by-parts, their sparse-matrix operators are closely related to multidimensional SBP operators [17]. The authors present two algorithms to solve for their sparse operators: a segregated method that solves for the diagonal mass matrix first and then solves for the coefficients of the stiffness matrices; and a coupled method that solves for the mass matrix and stiffness matrices simultaneously.
@@ 1.2 Paper outline
The bulk of the paper is devoted to describing our construction algorithm. Section 2 covers preliminary material and notation that we rely on in subsequent sections. In particular, Section 2 introduces the concept a degenerate SBP pair, in which the diagonal mass matrix is not necessarily positive definite. The construction of degenerate SBP pairs is the subject of Section 3, which begins with the construction of cell-based operators and then covers the assembly process for global operators. Section 4 explains how we convert a degenerate SBP pair into a diagonal-norm SBP operator by solving a linear inequality. The operators produced by the construction algorithm are studied and verified in Section 5. Section 6 concludes the paper with a summary and a discussion of potential future work.
@@ 2.1 Domain, background mesh, and nodes
We are interested in discretizing initial boundary-value problems that are posed on some bounded domain $\Omega\subset\mathbb{R}^{\mathsf{d}}$, where $\mathsf{d}$ is the spatial dimension. The boundary of $\Omega$ is denoted $\Gamma$, which may be partitioned into non-overlapping subsets, $\Gamma_{i}\subset\Gamma$, where distinct boundary conditions can be applied. An example domain is illustrated in Figure 1(a).
@@
We also need notation for the faces that make up the cell boundaries. Let $\{\Gamma^{\mathsf{f}}\}_{\mathsf{f}=1}^{N_{F}}$ denote the $N_{F}$ unique faces in the mesh. The subset of faces that coincide with the boundary of cell $\mathsf{c}$ will be denoted $F^{\mathsf{c}}\subset\{1,2,\ldots,N_{F}\}$. In addition, we will use $F_{I}\subset\{1,2,\ldots,N_{F}\}$ for the subset of all interfaces between cells, and $F_{B}\subset\{1,2,\ldots,N_{F}\}$ to denote the subset of faces that coincide with (or approximate) the boundary $\Gamma$.
@@ 2.2 Polynomial spaces and accuracy
The accuracy of various operators will be defined with respect to total degree polynomials; we will use $\mathbb{P}^{p}(\Omega)$ to indicate the space of total degree $p$ polynomials over an arbitrary domain $\Omega$.
We need a basis for $\mathbb{P}^{p}(\Omega^{\mathsf{c}})$ in order to define cell-based SBP operators. We will use the notation $\{\mathcal{V}_{j}\}_{j=1}^{N_{p,d}}$ when it is necessary to refer to the basis explicitly, where $N_{p,\mathsf{d}}={\scriptstyle\binom{p+\mathsf{d}}{\mathsf{d}}}$ is the dimension of the total degree $p$ basis in $\mathsf{d}$ dimensions. In practice, we adopt the Proriol orthogonal basis [23] for two-dimensional domains and their generalization for three-dimensional domains; see, e.g., [15, Chap. 10].
It is often necessary to evaluate the polynomial basis at some set of nodes. The Vandermonde matrix is the maxtrix $\mathsf{V}\in\mathbb{R}^{N\times N_{p,\mathsf{d}}}$ whose rows correspond to the nodes in $X$ and whose columns correspond to the basis vectors:
@@
Similarly, $\mathsf{V}^{\mathsf{c}}$ denotes the Vandermonde matrix of polynomials evaluated at the nodes $X^{\mathsf{c}}$ of cell $\mathsf{c}$.
@@ 2.3 First-derivative summation-by-parts operators
The above definition reframes first-derivative SBP operators in terms of $\mathsf{M}$ and $\mathsf{Q}_{x}$ rather than the difference operator $\mathsf{D}_{x}$; this allows us to consider operators with indefinite, or even singular, mass matrices. Otherwise, the degenerate SBP definition is essentially the same as the (diagonal-norm) multidimensional SBP definition introduced in [17]. Indeed, we can immediately relate diagonal-norm SBP operators to degenerate SBP pairs.
@@ 3.2.2 Stencil construction
The number of nodes in the stencil strongly influences the conditioning of the Vandermonde matrix, $\mathsf{V}^{\mathsf{c}}$, which then influences the properties of the degenerate SBP pair on cell $\mathsf{c}$. Therefore, similar to Kaur et al. [19], we add nodes to the stencil until the condition number of the degree $12p-$ 1 Vandermonde matrix — denoted $\mathsf{cond}(\mathsf{V}_{2p-1}^{\mathsf{c}})$ — is below some threshold $\mathsf{tol}$; see Algorithm 1. To generate the results for Section 5, we set the maximum number of inner iterations to $n_{\max}=4p-1$, and the condition-number threshold to $\mathsf{tol}=5\times 10^{2p-1}$.
A couple of points about Algorithm 1 are worth highlighting. First, the polynomial degree used in the algorithm is $12p-$ 1 rather than $p$, because the stencil must be sufficiently large to construct a $12p-$ 1 exact quadrature rule; this will be discussed more in Section 3.3.1. Second, since the nested loop in Algorithm 1 starts at $n=1$, the default stencil size is $N^{\mathsf{c}}=N_{2p-1,\mathsf{d}}+1$, which is one larger than the dimension of the total degree $12p-$ 1 polynomial basis. This default stencil size is motivated by our approach to ensure a positive norm described in Section 4.
@@ 3.3.1 Cell-based norm/quadrature
For the remainder of the Section 3, we will adopt the minimum-norm solution for the cell’s quadrature weights, $\bm{m}^{\mathsf{c}}=\bm{m}_{\min}^{\mathsf{c}}$. However, we will revisit this choice in Section 4, where the additional degrees of freedom provided by the null-space will be critical to ensuring a positive-definite mass matrix.
@@ 3.3.2 Cell-based symmetric part of $\mathsf{Q}_{x}^{\mathsf{c}}$
Based on the definitions of $\mathsf{R}^{\mathsf{f},\mathsf{c}}$, $\mathsf{B}^{\mathsf{f}}$ and $\mathsf{N}_{x}^{\mathsf{f},\mathsf{c}}$, it follows from [7, Theorem 1] that $\mathsf{E}_{x}^{\mathsf{c}}$ satisfies Equation (3) in Definition 1.
@@ 3.3.3 Cell-based skew-symmetric part of $\mathsf{Q}_{x}^{\mathsf{c}}$
Examining Table 1 suggests that some of the linear systems are over determined; however, the so-called compatibility conditions — see, e.g., [17, Lemma 3.1] — are embedded in these accuracy conditions. For example, for $p=1$, there are 6 redundant equations in two dimensions, and 10 redundant equations in three dimensions.
@@
Based on the accuracy of the quadrature rule in $\mathsf{M}$, and the fact that $\mathsf{E}_{x}$ satisfies Equation (3), the $ij$th entry in $\mathsf{V}_{x}\mathsf{M}\mathsf{V}-\frac{1}{2}\mathsf{V}^{T}\mathsf{E}_{x}\mathsf{V}$ is
@@ 3.4 Assembly of global SBP operators
Once the cell-based (degenerate) SBP pairs are constructed, they are assembled into global operators using an approach similar to [17, Theorem 4.1]. However, unlike the symmetric operators in [17], the $\mathsf{E}_{x}^{\mathsf{c}}$ in the present work are dense, so the assembly process must be modified to avoid symmetric contributions to $\mathsf{Q}_{x}$ from interior cells.
@@
Next, consider the degenerate SBP accuracy constraint, $\mathsf{Q}_{x}\mathsf{V}=\mathsf{M}\mathsf{V}_{x}$. Using the factorization $\mathsf{Q}_{x}=\mathsf{S}_{x}+\frac{1}{2}\mathsf{E}_{x}$ at both the global and cell levels, and substituting definitions (11) and (12), the left-hand side of the accuracy condition can be written as five distinct sums:
@@ 3.6 An operator for artificial dissipation
The matrices $\mathsf{R}_{\mathsf{diss}}^{\mathsf{f},+}$, $\mathsf{R}_{\mathsf{diss}}^{\mathsf{f},-}$, and $\mathsf{B}_{\mathsf{diss}}^{\mathsf{f}}$ are similar to the operators used in the construction of $\mathsf{S}_{x}$. The only difference is that the dissipation operators interpolate to only one point on the face — the geometric center — rather than all the face quadrature nodes. Consequently, the matrix $\mathsf{B}_{\mathsf{diss}}^{\mathsf{f}}$ is actually a scalar that holds the area of face $\mathsf{f}$.
@@ 4 Enforcing a positive-definite diagonal norm
Degenerate SBP operators with indefinite $\mathsf{M}$ can be used to solve steady-state problems, but their instability makes them inappropriate for unsteady problems. Unfortunately, for a given node distribution $X$, the mass matrix produced by Algorithm 2 is often indefinite for $p=1$ and almost always indefinite for $p\geq 2$.
@@
To the best of our knowledge, it is not possible to adjust the cell-based mass matrices independently in order to achieve a global, positive-definite mass matrix. This is because each node appears in the stencil of multiple cells, so changing $\mathsf{M}^{\mathsf{c}}$ to make one node’s (global) quadrature weight positive may cause another node’s weight to become negative.
@@ 4.2 When is the norm-inequality problem feasible?
We establish conditions for the feasibility of the norm-inequality problem by leveraging the theory of positive least-squares cubature formulas from [11]. One of the hypotheses required by the theory in [11] is that the nodes $X$ are part of an equidistributed sequence parameterized by $N$, e.g. Halton points [14]. Thus, in the statement of the theorem we use the notation $X_{N}$ to indicate that the node set includes all the points in the equidistributed sequence up to the $N$th.
@@
To conclude the proof using contradiction, suppose there are coefficients $\bm{\alpha}$ such that $\bm{u}=\mathsf{V}_{2p-1}\bm{\alpha}\geq\bm{0}$. Using the hypothesis that the nodes $X_{N}$ are equidistributed, we can apply [11, Corollary 3.6] to infer that there exists an $N^{*}$ such that, for all $N\geq N^{*}$, there is a least-squares cubature formula over $X_{N}$ with positive weights $\bm{w}\in\mathbb{R}^{N}$ that integrates polynomials in $\mathbb{P}^{2p-1}(\Omega)$ exactly. Using these weights with the assumption that $\bm{u}\geq\bm{0}$, we find
@@
The challenge is showing that the null-space is equal to the columns space of $\mathsf{V}_{2p-1}$. Proving this is beyond the scope of this work, but we offer some intuition regarding when it should hold. First, the row rank of $\mathsf{Z}$ must be $N-N_{2p-1,\mathsf{d}}$, otherwise there would another (nontrival) vector $\bm{u}\in\mathbb{R}^{N}$ that is not in the column space of $\mathsf{V}_{2p-1}$ but nevertheless satisfies $\bm{u}^{T}\mathsf{Z}=\bm{0}^{T}$. To ensure that $\mathsf{Z}$ has this maximal rank, its rows must be sufficiently “independent.” For example, if each node $i$ is in the stencil of a unique set of cells — i.e. no other node is associated with the same set of cells — then the sparsity pattern of each row of $\mathsf{Z}$ will be distinct. Unfortunately, while this might provide a path to proving the result we want, we hypothesize that it is a sufficient but not necessary condition.

### v2

@@ Abstract
High-order difference operators with the summation-by-parts (SBP) property can be used to build stable discretizations of hyperbolic conservation laws; however, most high-order SBP operators require a conforming, high-order mesh for the domain of interest. To circumvent this requirement, we present an algorithm for building high-order, diagonal-norm, first-derivative SBP operators on point clouds over level-set geometries. The algorithm is not mesh-free, since it uses a Cartesian cut-cell mesh to define the sparsity pattern of the operators and to provide intermediate quadrature rules; however, the mesh is generated automatically and can be discarded once the SBP operators have been constructed. Using this temporary mesh, we construct local, cell-based SBP difference operators that are assembled into global SBP operators. We identify conditions for the existence of a positive-definite diagonal mass matrix, and we compute the diagonal norm by solving a sparse system of linear inequalities using an interior-point algorithm. We also describe an artificial dissipation operator that complements the first-derivative operators when solving hyperbolic problems, although the dissipation is not required for stability. The numerical results confirm the conditions under which a diagonal norm exists and study the distribution of the norm’s entries. In addition, the results verify the accuracy and stability of the point-cloud SBP operators using the linear advection equation.
@@ 1 Introduction
The present contribution aims to accelerate both the run and pre-processing time of high-order SBP discretizations. Specifically, we present an algorithm to construct high-order, sparse SBP operators whose nodal degrees of freedom are defined on point clouds. The combination of these two elements — high-order SBP and point-cloud nodes — yields an accurate and stable discretization with considerable flexibility.
@@ 1.1 Related work
In [15], the authors present an algorithm to construct a high-order finite-difference method on point clouds. While they do not use the term summation-by-parts, their sparse-matrix operators are closely related to multidimensional SBP operators [12]. The authors present two algorithms to solve for their sparse operators: a segregated method that solves for the diagonal mass matrix first and then solves for the coefficients of the stiffness matrices; and a coupled method that solves for the mass matrix and stiffness matrices simultaneously. The coupled method has recently been extended to general function spaces by Glaubitz et al. [19]. In addition, Kwan and Chan [20] have proposed a simpler construction for a first-order variant of [15].
@@ 1.2 Paper outline
The bulk of the paper is devoted to describing our construction algorithm. Section 2 covers preliminary material and notation that we rely on in subsequent sections. In particular, Section 2 introduces the concept of a degenerate SBP pair, in which the diagonal mass matrix is not necessarily positive definite. The construction of degenerate SBP pairs is the subject of Section 3, which begins with the construction of cell-based operators and then covers the assembly process for global operators. Section 4 explains how we convert a degenerate SBP pair into a diagonal-norm SBP operator by solving a linear inequality. The operators produced by the construction algorithm are studied and verified in Section 5. Section 6 concludes the paper with a summary and a discussion of potential future work.
@@ 2.1 Domain, background mesh, and nodes
We are interested in discretizing initial boundary-value problems that are posed on some bounded domain $\Omega\subset\mathbb{R}^{\mathsf{d}}$, where $\mathsf{d}$ is the spatial dimension. The boundary of $\Omega$ is denoted $\Gamma$, which may be partitioned into non-overlapping subsets, $\Gamma_{i}\subset\Gamma$, where distinct boundary conditions can be applied. An example domain — a unit box with a circle of radius $\frac{1}{4}$ excised from its center — is illustrated in Figure 1(a).
@@
We also need notation for the faces that make up the cell boundaries. Let $\{\Gamma^{\mathsf{f}}\}_{\mathsf{f}=1}^{N_{F}}$ denote the $N_{F}$ unique faces in the mesh. The subset of faces that coincide with the boundary of cell $\mathsf{c}$ will be denoted $F^{\mathsf{c}}\subset\{1,2,\ldots,N_{F}\}$. In addition, we will use $F_{I}\subset\{1,2,\ldots,N_{F}\}$ for the subset of all interfaces between cells, and $F_{B}\subset\{1,2,\ldots,N_{F}\}$ to denote the subset of faces that coincide with (or approximate) the boundary $\Gamma$. In addition to Figure 1(b), the various face types are illustrated in Figure 2.
@@ 2.2 Polynomial spaces and accuracy
The accuracy of various operators will be defined with respect to total degree polynomials. While the accuracy of SBP operators can be defined using other sets of functions [13, 18], we leave such an extension to future work. We will use $\mathbb{P}^{p}(\Omega)$ to indicate the space of total degree $p$ polynomials over an arbitrary domain $\Omega$.
We need a basis for $\mathbb{P}^{p}(\Omega^{\mathsf{c}})$ in order to define cell-based SBP operators. We will use the notation $\{\mathcal{V}_{j}\}_{j=1}^{N_{p,d}}$ when it is necessary to refer to the basis explicitly, where $N_{p,\mathsf{d}}={\scriptstyle\binom{p+\mathsf{d}}{\mathsf{d}}}$ is the dimension of the total degree $p$ basis in $\mathsf{d}$ dimensions. In practice, we adopt the Proriol orthogonal basis [22] for two-dimensional problems and their generalization for three-dimensional problems — see, e.g., [23, Chap. 10] — although other choices are certainly possible.
It is often necessary to evaluate the polynomial basis at some set of nodes. The Vandermonde matrix is the matrix $\mathsf{V}\in\mathbb{R}^{N\times N_{p,\mathsf{d}}}$ whose rows correspond to the nodes in $X$ and whose columns correspond to the basis vectors:
@@
Similarly, $\mathsf{V}^{\mathsf{c}}$ denotes the Vandermonde matrix of polynomials evaluated at the nodes $X^{\mathsf{c}}$ of cell $\mathsf{c}$. In practice, we evaluate the Vandermonde matrix by first translating and scaling the nodes so that they lie within the reference domain where the basis is orthogonal, i.e., the reference triangle $\{x,y>-1,x+y<0\}$, in two dimensions, and the reference tetrahedron $\{x,y,z>-1,x+y+z<-1\}$, in three dimensions. Transforming the nodes to lie within the reference domain tends to improve the conditioning of the Vandermonde matrix.
@@ 2.3 First-derivative summation-by-parts operators
The above definition reframes first-derivative SBP operators in terms of $\mathsf{M}$ and $\mathsf{Q}_{x}$ rather than the difference operator $\mathsf{D}_{x}$; this allows us to consider operators with indefinite, or even singular, mass matrices $\mathsf{M}$. Otherwise, the degenerate SBP definition is essentially the same as the (diagonal-norm) multidimensional SBP definition introduced in [12]. Indeed, we can immediately relate diagonal-norm SBP operators to degenerate SBP pairs.
@@ 3.2 Background mesh
Remark 2.
We limit each cell to one node because this condition is easy to enforce, but it is not an inherent requirement of the method and one can have multiple nodes per cell.
@@ 3.2.2 Stencil construction
The number of nodes in the stencil strongly influences the conditioning of the Vandermonde matrix, $\mathsf{V}^{\mathsf{c}}$, which then influences the properties of the degenerate SBP pair on cell $\mathsf{c}$. Therefore, similar to Kaur et al. [17], we add nodes to the stencil until the condition number of the degree $12p-$ 1 Vandermonde matrix — denoted $\mathsf{cond}(\mathsf{V}_{2p-1}^{\mathsf{c}})$ — is below some threshold $\mathsf{tol}$ or a maximum number of iterations $n_{\max}$ is reached; see Algorithm 1. To generate the results for Section 5, we set the maximum number of inner iterations to $n_{\max}=4p-1$, and the condition-number threshold to $\mathsf{tol}=5\times 10^{2p-1}$. In our numerical experiments, the innermost loop of Algorithm 1 always terminated before reaching $n_{\max}$; nevertheless, adding nodes is not guaranteed to reduce the condition number of the Vandermonde matrix, and some node sets and geometries may require larger $n_{\max}$ or even node repositioning.
A couple of points about Algorithm 1 are worth highlighting. First, the polynomial degree used in the algorithm is $12p-$ 1 rather than $p$, because the stencil must be sufficiently large to construct a $12p-$ 1 exact quadrature rule; this will be discussed more in Section 3.3.1. Second, since the nested loop in Algorithm 1 starts at $n=1$, the default stencil size is $N^{\mathsf{c}}=N_{2p-1,\mathsf{d}}+1$, which is one larger than the dimension of the total degree $12p-$ 1 polynomial basis. This default stencil size is motivated by our approach to ensure a positive norm. In short, the larger-than-necessary stencil introduces additional degrees of freedom in the cell-based quadrature that can be used to enforce positivity; see Section 4.
@@ 3.3.1 Cell-based norm/quadrature
For the remainder of Section 3, we will adopt the minimum-norm solution for the cell’s quadrature weights, $\bm{m}^{\mathsf{c}}=\bm{m}_{\min}^{\mathsf{c}}$. However, we will revisit this choice in Section 4, where the additional degrees of freedom provided by the null-space will be critical to ensuring a positive-definite mass matrix.
@@ 3.3.2 Cell-based symmetric part of $\mathsf{Q}_{x}^{\mathsf{c}}$
Based on the definitions of $\mathsf{R}^{\mathsf{f},\mathsf{c}}$, $\mathsf{B}^{\mathsf{f}}$ and $\mathsf{N}_{x}^{\mathsf{f},\mathsf{c}}$, it follows from [27, Theorem 1] that $\mathsf{E}_{x}^{\mathsf{c}}$ satisfies Equation (3) in Definition 1.
@@ 3.3.3 Cell-based skew-symmetric part of $\mathsf{Q}_{x}^{\mathsf{c}}$
Examining Table 1 suggests that some of the linear systems are over determined; however, the so-called compatibility conditions — see, e.g., [12, Lemma 3.1] — are embedded in these accuracy conditions. For example, for $p=1$, there are 6 redundant equations in two dimensions, and 10 redundant equations in three dimensions.
@@
Based on the accuracy of the quadrature rule in $\mathsf{M}$, and the fact that $\mathsf{E}_{x}$ satisfies Equation (3), the $ij$th entry in $\mathsf{V}_{x}^{T}\mathsf{M}\mathsf{V}-\frac{1}{2}\mathsf{V}^{T}\mathsf{E}_{x}\mathsf{V}$ is
@@ 3.4 Assembly of global SBP operators
Once the cell-based (degenerate) SBP pairs are constructed, they are assembled into global operators using an approach similar to [12, Theorem 4.1]. However, unlike the symmetric operators in [12], the $\mathsf{E}_{x}^{\mathsf{c}}$ in the present work are dense, so the assembly process must be modified to avoid symmetric contributions to $\mathsf{Q}_{x}$ from interior cells.
@@
Next, consider the degenerate SBP accuracy constraint, $\mathsf{Q}_{x}\mathsf{V}=\mathsf{M}\mathsf{V}_{x}$. Using the factorization $\mathsf{Q}_{x}=\mathsf{S}_{x}+\frac{1}{2}\mathsf{E}_{x}$ at both the global and cell levels, and substituting definitions (11) and (12), the left-hand side of the accuracy condition (1) can be written as five distinct sums (see A for more details):
@@ 3.6 An operator for artificial dissipation
The matrices $\mathsf{R}_{\mathsf{diss}}^{\mathsf{f},+}$, $\mathsf{R}_{\mathsf{diss}}^{\mathsf{f},-}$, and $\mathsf{B}_{\mathsf{diss}}^{\mathsf{f}}$ are similar to the operators used in the construction of $\mathsf{S}_{x}$. The only difference is that the dissipation operators interpolate to only one point on the face — the geometric center — rather than all the face quadrature nodes. Consequently, the matrix $\mathsf{B}_{\mathsf{diss}}^{\mathsf{f}}$ is actually a scalar that holds the area of face $\mathsf{f}$. Limiting the interpolation to one point on the face reduces the computational cost of applying the dissipation operator, which would otherwise grow in proportion to the number of face quadrature points, i.e. $(p+1)^{\mathsf{d}}$ for the Gauss-Legendre rules used here.
@@ 4 Enforcing a positive-definite diagonal norm
Degenerate SBP operators with indefinite $\mathsf{M}$ can be used to solve steady-state problems, because the weak formulation of the discretization only requires the diagonal norm to be a sufficiently accurate quadrature. On the other hand, an indefinite $\mathsf{M}$ is not suitable for unsteady problems since the discretization will be unstable, in general. Unfortunately, for a given node distribution $X$, the mass matrix produced by Algorithm 2 is often indefinite for $p=1$ and almost always indefinite for $p\geq 2$.
@@
To the best of our knowledge, it is not possible to adjust the cell-based mass matrices independently in order to achieve a global, positive-definite mass matrix. This is because the diagonal entries of $\mathsf{M}^{\mathsf{c}}$ are coupled to one another — see Equation (6) — such that increasing the value of one entry may decrease another entry. Consequently, when attempting to make one node’s (global) quadrature weight positive we may inadvertently make another weight negative.
@@ 4.2 When is the norm-inequality problem feasible?
[…]
@@
To conclude the proof using contradiction, suppose there are coefficients $\bm{\alpha}$ such that $\bm{u}=\mathsf{V}_{2p-1}\bm{\alpha}\geq\bm{0}$. Using the hypothesis that the nodes $X_{N}$ are equidistributed, we can apply [28, Corollary 3.6] to infer that there exists an $N^{*}$ such that, for all $N\geq N^{*}$, there is a least-squares cubature formula over $X_{N}$ with positive weights $\bm{w}\in\mathbb{R}^{N}$ that integrates polynomials in $\mathbb{P}^{2p-1}(\Omega)$ exactly. Using these weights with the assumption that $\bm{u}\geq\bm{0}$, we find
@@
The challenge is showing that the null-space is equal to the column space of $\mathsf{V}_{2p-1}$. Proving this is beyond the scope of this work, but we offer some intuition regarding when it should hold. First, the row rank of $\mathsf{Z}$ must be $N-N_{2p-1,\mathsf{d}}$, otherwise there would another (nontrival) vector $\bm{u}\in\mathbb{R}^{N}$ that is not in the column space of $\mathsf{V}_{2p-1}$ but nevertheless satisfies $\bm{u}^{T}\mathsf{Z}=\bm{0}^{T}$. To ensure that $\mathsf{Z}$ has this maximal rank, its rows must be sufficiently “independent.” For example, if each node $i$ is in the stencil of a unique set of cells — i.e. no other node is associated with the same set of cells — then the sparsity pattern of each row of $\mathsf{Z}$ will be distinct. Unfortunately, while this might provide a path to proving the result we want, we hypothesize that it is a sufficient but not necessary condition.